OpenStudy (anonymous):

apply the quadratic formula to find the roots of the given function, and then graph the function. f(x) = y2 - 9 show work please

OpenStudy (curry):

(y+3)(y-3) so y=3 and y=-3 which are the roots

OpenStudy (curry):

do u know the formula for the quadratic equation

OpenStudy (anonymous):

Yes, \[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

not sure how \[f(x)=y^2-9\] makes sense

OpenStudy (anonymous):

if it is \[f(x)=x^2-9\] then the solutions to \[x^2-9=0\] are 3 and -3 with your eyeballs, not with any formula \[x^2-9=0\] \[x^2-9\] \[x=\pm3\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

yw